A grating has 764 lines per centimeter.

1) Find the angles of the first three principal maxima above the central fringe when this grating is illuminated with 643 nmnm light.

Enter your answers numerically separated by commas.
Answer #1

# Shit width d= 16 1 764%.01 d=1.309+10-6 m * first maxima ?m=13 dsino = ma >> 1-309310she = 16434169 O = 29.4230 & Ind maxim